1. Let S P(ta), P(ld) if you are given that {a,b,c,d) be a sample space and A-{a,b,e). В-{a,b,d) and C-{b,e). Calculate the probabilities a) P(A) = , P(B)-i and P(C)- 21 b) P(C)P(A) = P(A)-P(C)-흄
Answer #1

1)a) P(a) =P(A)-P(C)=1/2-10/21 =1/42

P(d) =P(S)-P(A)=1-1/2 =1/2

b)here P(C)*P(A)=1/3........(1)

P(A)-P(C')=P(A)-(1-P(C))=P(A)+P(C)-1=1/3

P(A)+P(C)=4/3 .......(2)

solving equation (1) and (2)

P(A)=1

and P(C)=1/3

therefore

P(a) =P(A)-P(C)=1-1/3 =2/3

P(d) =P(S)-P(A)=1-1 =0